OpenStudy (anonymous):

Lars deposited $50 into a savings account for which interest is compounded quarterly. According to the rule of 72, what interest rate will cause his money to double in approximately 29 years?

OpenStudy (anonymous):

A. 0.6% B. 1.7% C. 2.5% D. 0.4%

OpenStudy (wolf1728):

Total = princ * (1+rate)^n let's say when will $100 become $200 log(1 + rate) = {log(total) -log(Principal)} ÷ Years log(1 + rate) = [log (200) - log(100)] / Years log(1 + rate) = 2.3010299957 - 2 / 29 log(1 + rate) = .3010299957 / 29 log (1 + rate) = 0.0103803447 1 + rate = 10^0.0103803447 1 + rate = 1.0241895602 rate = .0241895602 or rate = 2.41895602 % Then the answer is C

OpenStudy (wolf1728):

I didn't use the rule of 72 and I used a method that gives an exact answer.

OpenStudy (wolf1728):

Okay to use the rule of 72: time to double = 72 / rate rate = 72 / time to double rate = 72 / 29 = 2.4827586207 rate approximately equals 2.5%